下面的茎叶图显示了30名学生在满分为80分的科学测试中所获得的分数。

  1. (i)

    求中位数分数。

    答案 ...

  2. (ii)

    解释为什么平均数可能不是用来概括这些学生所获分数的合适代表值。

    答案 ...

    ...

  3. (iii)

    成绩前30%的学生可获得优秀。

    求学生获得优秀所需的最低分数。

    答案 ...

  4. (iv)

    求分数的标准差。

    答案 ...

  5. (v)

    第二组参加同一测试的学生所获分数的标准差为9.85。

    利用这一信息,对两组学生所获分数进行评价。

    答案 ...

    ...

  6. (8)
    1. (b)

      在商店的收银台,顾客可以使用NETS(电子支付)、信用卡或现金进行付款。

      顾客使用NETS付款的概率为 23\tfrac{2}{3},使用信用卡付款的概率为 14\tfrac{1}{4}。

      1. (i)

        求顾客用现金付款的概率。

        答案 ...

      2. (ii)

        两名顾客正在为他们购买的商品付款。

        求以下事件的概率:

        1. (a)

          两名顾客都使用信用卡付款,

          答案 ...

        2. (b)

          他们在付款时未使用现金。

          答案 ...

      3. (iii)

        三名顾客正在为他们购买的商品付款。

        求第一次现金交易发生在第三位顾客(而非之前)的概率。

        答案 ...

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

8 (a)(i)

1. Find the positions of the two middle marks

There are 3030 students.

For an even number of values, the median is the average of the 302=15\frac{30}{2}=15th and 1616th values.

2. Identify the 15th and 16th values from the ordered stem-and-leaf diagram and average them

Counting from the smallest mark:

1818 is the 1st value.

The 40s give the 2nd to 9th values.

The 50s are 50,51,51,52,54,58,5950,51,51,52,54,58,59, which are the 10th to 16th values.

So the 15th value is 5858 and the 16th value is 5959.

Median=58+592=58.5\text{Median}=\frac{58+59}{2}=58.5

答案

58.558.5 marks

8 (a)(ii)

1. Explain why the mean is not a suitable average for this data set

From the stem-and-leaf diagram, most marks are from the 4040s to 7070s, but there is one very low mark, 1818. This outlier would pull the mean down, so the mean may not be representative of the marks obtained by the students.

答案

The mean may not be appropriate because there is an outlier, 1818, which is much lower than the other marks, so it would lower the mean and not represent the students’ marks well.

8 (a)(iii)

1. Find how many students are awarded distinction and identify the cut-off position

Top 30%30\% of 3030 students:
0.30×30=90.30 \times 30 = 9
So the top 99 students get distinction.

In ascending order, these are the 2222nd to 3030th marks, since
30−9+1=22.30-9+1=22.

2. Read the 22nd mark from the stem-and-leaf diagram

Counting from the lowest mark:

  • 1818 is the 1st mark,
  • the 4040-marks are the 2nd to 9th marks,
  • the 5050-marks are the 10th to 16th marks,
  • the 6060-marks are the 17th to 23rd marks:

63, 63, 63, 63, 66, 67, 6763,\ 63,\ 63,\ 63,\ 66,\ 67,\ 67

Hence the 2222nd mark is 6767.

答案

6767 marks

8 (a)(8)(b)(i)

1. Subtract the probabilities of NETS and credit card from 1

P(cash)=1−P(NETS)−P(credit card)P(\text{cash})=1-P(\text{NETS})-P(\text{credit card})
=1−23−14=1-\frac{2}{3}-\frac{1}{4}
=1212−812−312=112=\frac{12}{12}-\frac{8}{12}-\frac{3}{12}=\frac{1}{12}

答案

112\dfrac{1}{12}

8 (a)(8)(b)(ii)(a)

1. Find the probability that both customers pay by credit card

For one customer,
P(credit card)=14P(\text{credit card})=\dfrac{1}{4}.

For two customers,
P(both credit card)=14×14=116.P(\text{both credit card})=\dfrac{1}{4}\times \dfrac{1}{4}=\dfrac{1}{16}.

答案

116\dfrac{1}{16}

8 (a)(8)(b)(ii)(b)

1. Find the probability that one customer does not pay by cash

From part (b)(i),
P(cash)=112P(\text{cash})=\dfrac{1}{12}.

So
P(no cash)=1−112=1112P(\text{no cash})=1-\dfrac{1}{12}=\dfrac{11}{12}.

2. Find the probability that both customers do not use cash

For two customers,
P(no cash for both)=1112×1112=121144.P(\text{no cash for both})=\dfrac{11}{12}\times\dfrac{11}{12}=\dfrac{121}{144}.

答案

121144\dfrac{121}{144}

8 (a)(8)(b)(iii)

1. Identify the probability of not paying by cash and paying by cash

From the earlier part,
P(cash)=112.P(\text{cash})=\frac{1}{12}.
So
P(not cash)=1−112=1112.P(\text{not cash})=1-\frac{1}{12}=\frac{11}{12}.

2. Multiply the probabilities for the sequence not cash, not cash, cash

For the first cash transaction to occur with the third customer, the first two customers must not pay by cash and the third customer must pay by cash.
P=\frac{11}{12}\times\frac{11}{12}\times\frac{1}{12}= rac{121}{1728}.

答案

1211728\dfrac{121}{1728}

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1某班第二组女生参加体育测试,仰卧起坐的成绩(单位:个)如下:43、41、39、40、37.这组数据的中位数是 .
  2. 2车间生产同一种产品,某一星期日产量的情况如下:有22天是154154件,有33天是152152件,有22天是146146件,那么车间这个星期日产量的众数是 ___.
  3. 3茎叶图显示了某班男生和女生的测试成绩。 [茎叶图印刷为两列,左列标为“男生”(Boys),右列标为“女生”(Girls),共用中间的一列茎。键(男生)8 | 3 表示 38。键(女生)4 | 1 表示 41。] 艾尔文在下方的箱线图中表示了…
  4. 4有三个数5,x,9,它们的平均数为6,则x为 .
  5. 5在一组由五个数字组成的集合中,最小的数字是一个偶数,且它是质数。极差和平均数都是 7。中位数也是唯一的众数。写出这五个数字。